# SLIDE: A Framework Integrating Small and Large Language Models for Open-Domain Dialogues Evaluation

### v1

@@ Abstract
Kun Zhao1thanks:  Equal contribution. , Bohao Yang2footnotemark: , Chen Tang2, Chenghua Lin2thanks:  Corresponding authors. , Liang Zhan1footnotemark: 1 Department of Electrical and Computer Engineering, University of Pittsburgh, US 2 Department of Computer Science, The University of Manchester, UK {kun.zhao, liang.zhan}@pitt.edu, bohao.yang-2@postgrad.manchester.ac.uk {chen.tang, chenghua.lin}@manchester.ac.uk
@@ 1 Introduction
Open-domain dialogue generation is an important research topic in the field of Natural Language Generation (NLG) Zhao et al. (2023); Xiao et al. (2023); Ye et al. (2024); Tang et al. (2023b); Loakman et al. (2023); Yang et al. (2023); Tang et al. (2023a); Dai et al. (2024); Yang et al. (2024a). Evaluating such dialogues, however, is challenging due to the prevalent one-to-many issue where one conversational context may have multiple reasonable, yet semantically different responses. Additionally, adversarial negative responses, which have word overlap with given contexts, also raise further challenges in adequately evaluating open-domain dialogue Sai et al. (2020).
Existing evaluation methods centered on word-overlap, such as BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005)) and semantic-embedding based approaches (e.g., BERTScore Zhang et al. (2020) and BARTScore Yuan et al. (2021)), evaluate responses via calculating the similarity between the generated response and contexts or gold references. Therefore, they struggle to adequately address the one-to-many problem as a context may permit many different responses. Recently, there are some novel approaches proposed to leverage the commonsense reasoning capabilities learned in Large Language Models (LLMs) for automatic evaluation Fu et al. (2023); Liu et al. (2023); Kocmi and Federmann (2023); Chiang and yi Lee (2023); Yang et al. (2024b); Tang et al. (2024). However, there are still some limitations of LLM-based evaluation. For example, Lin and Chen (2023) propose LLM-EVAL to evaluate open-domain dialogue systems with a range of LLMs and identify numerous issues including the significant reliance on prompt phrasing, which may artificially limit the effectiveness of the metrics and the evaluation quality changed with prompt formulation, which makes scoring unreliable and unrobust. Additionally, Wang et al. (2023b) state that without additional support, LLMs alone do not constitute equitable evaluators.
@@ 2.2 LLM-based Evaluators
The application of LLM for evaluation purposes has seen significant interest due to its impressive capabilities across various tasks. GPTScore was introduced by Fu et al. (2023) as a method for multi-faceted, tailor-made, and training-independent evaluation. Preliminary investigations by Wang et al. (2023a) into the efficacy of evaluators based on LLM have been conducted. Kocmi and Federmann (2023) have utilised GPT models for assessing machine translation quality. Furthermore, Liu et al. (2023) developed GPT-EVAL, leveraging GPT-4 to evaluate across multiple tasks such as dialogue response generation, text summarization, data-to-text generation, and machine translation. Despite these advancements, the use of LLM-based metrics to assess adversarial negative responses within open-domain dialogue scenarios remains unexplored.
@@ 3.3 Evaluation Process
After training the SLM, we proceed to the evaluation phase. We first follow G-EVAL Liu et al. (2023) and prompt LLM to evaluate the given context and response. The prompt could be found in  A.1. Therefore we obtain a score referred to as $\mathrm{Score_{LLM}}$. Subsequently, SLM is employed to encode both the context and the response. This step involves segregating the response encoding into two distinct embeddings—robust and non-robust. The mathematical representation of this process is given by
@@ 4.2 Baselines
We select a range of widely used baseline metrics. The word-ovelap and embedding-based metrics includs BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005), Embedding-Average Wieting et al. (2016), Vector-Extrema Forgues et al. (2014), BERTScore Zhang et al. (2020), Unieval Zhong et al. (2022), and BARTScore Yuan et al. (2021). For the LLM-based metrics, we select recently proposed baselines, including G-EVAL Liu et al. (2023) and another LLM-based metric proposed by Chiang and yi Lee (2023) which we denote LLM-Chiang, as well as Gemini.
@@ 4.4.1 Dialogue Classification Task
As Table 1 shows, SLM employs both probability and distance measures, achieves superior performance with an accuracy of 91.05%, which exceeds the 88.91% accuracy achieved by GPT-4. Experimental results show that the SLM particularly excels in the classification of positive responses, achieving the highest accuracy of 91.83%. Conversely, GPT-4 exhibits the strongest performance in classifying adversarial negative responses, achieving 97.40% accuracy. The performance of Gemini surpasses that of GPT-3.5, yet remains inferior to that of GPT-4. These results indicate that while LLMs, such as GPT-4, possess robust capabilities in recognizing adversarial negative responses, they still struggle with the one-to-many problem, especially when dealing with semantically diverse positive responses. On the other hand, SLM with the integration of distance and probability metrics shows exceptional performance in recognizing open-domain positive responses. Based on these insights, we propose a hybrid approach that combines the SLM, a task-specific model, with LLMs for comprehensive dialogue evaluation. This approach is designed to leverage the distinct strengths of both types of models to enhance the overall performance in open-domain dialogevaluations.
[b] DailyDialog++ PersonaChat TopicalChat Metrics Pearson’s $\rho$ Spearman’s $\tau$ Pearson’s $\rho$ Spearman’s $\tau$ Pearson’s $\rho$ Spearman’s $\tau$ BLEU-1 0.303 (<0.0001) 0.242 (<0.0001) 0.331 (<0.0001) 0.276 (<0.0001) 0.352 (<0.0001) 0.330 (<0.0001) BLEU-2 0.317 (<0.0001) 0.270 (<0.0001) 0.311 (<0.0001) 0.261 (<0.0001) 0.279 (<0.0001) 0.252 (<0.0001) BLEU-3 0.256 (<0.0001) 0.251 (<0.0001) 0.251 (<0.0001) 0.239 (<0.0001) 0.179 (<0.0001) 0.189 (<0.0001) BLEU-4 0.234 (<0.0001) 0.247 (<0.0001) 0.214 (<0.0001) 0.217 (<0.0001) 0.247 (<0.0001) 0.154 (<0.0001) ROUGE-1 0.303 (<0.0001) 0.270 (<0.0001) 0.304 (<0.0001) 0.250 (<0.0001) 0.374 (<0.0001) 0.356 (<0.0001) ROUGE-2 0.178 (<0.0001) 0.133 (<0.0001) 0.142 (<0.0001) 0.126 (<0.0001) 0.173 (<0.0001) 0.149 (<0.0001) ROUGE-L 0.305 (<0.0001) 0.271 (<0.0001) 0.294 (<0.0001) 0.244 (<0.0001) 0.279 (<0.0001) 0.259 (<0.0001) METEOR 0.196 (<0.0001) 0.142 (<0.0001) 0.250 (<0.0001) 0.200 (<0.0001) 0.293 (<0.0001) 0.278 (<0.0001) Embedding Extrema 0.382 (<0.0001) 0.367 (<0.0001) 0.330 (<0.0001) 0.299 (<0.0001) 0.389 (<0.0001) 0.244 (<0.0001) Greedy 0.322 (<0.0001) 0.272 (<0.0001) 0.328 (<0.0001) 0.290 (<0.0001) 0.272(<0.0001) 0.248 (<0.0001) Average 0.186 (<0.0001) 0.208 (<0.0001) 0.239 (<0.0001) 0.249 (<0.0001) 0.254(<0.0001) 0.261 (<0.0001) BERTScore 0.419 (<0.0001) 0.381 (<0.0001) 0.465 (<0.0001) 0.412 (<0.0001) 0.525 (<0.0001) 0.485 (<0.0001) BARTSCORE 0.305 (<0.0001) 0.264 (<0.0001) 0.466 (<0.0001) 0.442 (<0.0001) 0.470 (<0.0001) 0.427 (<0.0001) Unieval 0.103 (<0.0001) 0.069 (<0.0001) 0.192 (<0.0001) 0.131 (<0.0001) 0.357 (<0.0001) 0.290 (<0.0001) G-EVAL (GPT-3.5) 0.447 (<0.0001) 0.414 (<0.0001) 0.516 (<0.0001) 0.562 (<0.0001) 0.642 (<0.0001) 0.668 (<0.0001) G-EVAL(GPT-4) 0.641 (<0.0001) 0.570 (<0.0001) 0.597 (<0.0000) 0.663 (<0.0001) 0.696 (<0.0001) 0.729 (<0.0001) LLM-Chiang (GPT-3.5) 0.632 (<0.0001) 0.583 (<0.0001) 0.599 (<0.0001) 0.670 (<0.0001) 0.667 (<0.0001) 0.661 (<0.0001) LLM-Chiang (GPT-4) 0.723 (<0.0001) 0.701 (<0.0001) 0.701(<0.0001) 0.675 (<0.0001) 0.678 (<0.0001) 0.709 (<0.0001) Ours(SLM-Prob-only) 0.677 (<0.0001) 0.632 (<0.0001) 0.638 (<0.0001) 0.692 (<0.0001) 0.691 (<0.0001) 0.689 (<0.0001) Ours(SLM-Dis-only) 0.652 (<0.0001) 0.600 (<0.0001) 0.668 (<0.0001) 0.692 (<0.0001) 0.642 (<0.0001) 0.673 (<0.0001) Ours(SLM-Dis and Prob) 0.728 (<0.0001) 0.666 (<0.0001) 0.711 (<0.0001) 0.707 (<0.0001) 0.741 (<0.0001) 0.727(<0.0001) Ours(LLMs-only (GPT-3.5)) 0.518 (<0.0001) 0.472 (<0.0001) 0.506 (<0.0001) 0.617 (<0.0001) 0.590 (<0.0001) 0.669 (<0.0001) SLIDE(GPT-3.5)) 0.727 (<0.0001) 0.674 (<0.0001) 0.692 (<0.0001) 0.695 (<0.0001) 0.696 (<0.0001) 0.718 (<0.0001) Ours(LLMs-only (Gemini)) 0.6600 (<0.0001) 0.6140 (<0.0001) 0.7199 (<0.0001) 0.6882 (<0.0001) 0.7015 (<0.0001) 0.6993(<0.0001) SLIDE(Gemini) 0.7600 (<0.0001) 0.6890 (<0.0001) 0.7011 (<0.0001) 0.7052 (<0.0001) 0.7138 (<0.0001) 0.7121 (<0.0001) Ours(LLMs-only (GPT-4)) 0.709 (<0.0001) 0.666 (<0.0001) 0.737 (<0.0001) 0.729 (<0.0001) 0.745 (<0.0001) 0.736 (<0.0001) SLIDE(GPT-4) 0.773 (<0.0001) 0.704 (<0.0001) 0.713 (<0.0001) 0.709 (<0.0001) 0.711 (<0.0001) 0.722 (<0.0001)
@@ 4.4.2 Dialogue Evaluation
Figure 2: T-SNE visualisation of the sentence representation of context and responses. The left panel, labeled "Normal," illustrates the vectors prior to disentanglement, whereas the right panel, labeled "Disentangled," displays the post-disentanglement outcomes. This demonstrates the convergence of negative responses towards the context following disentanglement.
4.5 Example Visualization
To demonstrate the effectiveness of disentanglement, we selected a subset of examples from the DailyDialog++ dataset and employed T-SNE for visualization, as depicted in Figure 2. Each example includes a context, five positive responses, and five adversarial negative responses. The vector representations prior to disentanglement are illustrated in the left plot, whereas those post-disentanglement are shown on the right. Post-disentanglement visualizations reveal a more distinct clustering of responses in relation to their contexts. By utilising these refined representations to train a classifier, we observed an improvement in classification accuracy on the test set of DailyDialog++ dataset, from 87% to 88.19%. Additional examples are available in Appendix A.5.
4.6 Case Study
5 Conclusion
@@
FS: We have another traditional holiday-the Dragon Boat Festival. SS: When is it? FS: It falls on the fifth day of the fifth lunar month.
@@
FS: We have been over this a hundred times! We are not getting a pet! SS: Why not? Come on! Just a cute little puppy or a kitty! FS: Who is going to look after a dog or a cat?
@@
2.0 2.79 2.4
Table 3: Samples from the DailyDialog++ dataset.
@@ A.4 Case Study
FS: Of course . Let me tell you in some detail about our idea.You know the popular Hello Kitty products. SS: Yes, of course. FS: Well, the products in themselves are very simple.It is the logo that is successful.So, Hello Kitty is successful because of the logo, but the products are very simple.
Reference:
Well, the brand name plays an important role in business.
Response:
Yeah, I don’t think so, the products must be good.
Human G-EVAL(GPT-3.5) G-EVAL(GPT-4)
4 2.25 2
Chiang(GPT-3.5) Chiang(GPT-4) SLM
2.5 1.75 4.6
ours(w/o LLM) SLIDE(GPT-3.5) SLIDE(GPT-4)
5 4.4 4
Context:
FS: We have been over this a hundred times! We are not getting a pet! SS: Why not? Come on! Just a cute little puppy or a kitty! FS: Who is going to look after a dog or a cat?
@@
FS: Why not? We’re supposed to meet him there. SS: Why doesn’t he meet us outside? FS: Why should he do that? It isn’t illegal for us to go in.
@@
5 5 4.8
Table 4: Samples from the DailyDialog++ dataset.
A.5 Example Visualization

### v2

@@ Abstract
Kun Zhao1thanks:  Equal contribution , Bohao Yang2footnotemark: , Chen Tang2, Chenghua Lin2thanks:  Corresponding authors , Liang Zhan1footnotemark: 1 Department of Electrical and Computer Engineering, University of Pittsburgh, US 2 Department of Computer Science, The University of Manchester, UK {kun.zhao, liang.zhan}@pitt.edu, bohao.yang-2@postgrad.manchester.ac.uk {chen.tang, chenghua.lin}@manchester.ac.uk
@@ 1 Introduction
Open-domain dialogue generation is an important research topic in the field of Natural Language Processing (NLP) Zeng et al. (2021); Wang et al. (2021); Xiao et al. (2023); Tang et al. (2023b, a); Yang et al. (2024a). Evaluating such dialogues, however, is challenging due to the prevalent one-to-many issue where one conversational context may have multiple reasonable, yet semantically different responses. Additionally, adversarial negative responses, which have word overlap with given contexts, also raise further challenges in adequately evaluating open-domain dialogue Sai et al. (2020).
Existing evaluation methods centered on word-overlap, such as BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005)) and semantic-embedding based approaches (e.g., BERTScore Zhang et al. (2020) and BARTScore Yuan et al. (2021)), evaluate responses via calculating the similarity between the generated response and contexts or gold references. Therefore, they struggle to adequately address the one-to-many problem as a context may permit many different responses. Recently, there are some novel approaches proposed to leverage the commonsense reasoning capabilities learned in Large Language Models (LLMs) for a wide range of NLP tasks, such as dialogue generation, evaluation and sentiment analysis Loakman et al. (2023); liu-etal-2023-g; Yang et al. (2024b); Chiang and yi Lee (2023); Yang et al. (2024c). However, there are still some limitations of LLM-based evaluation. For example, Lin and Chen (2023) propose LLM-EVAL to evaluate open-domain dialogue systems with a range of LLMs and identify numerous issues including the significant reliance on prompt phrasing, which may artificially limit the effectiveness of the metrics and the evaluation quality changed with prompt formulation, which makes scoring unreliable and unrobust. Additionally, Wang et al. (2023b) state that without additional support, LLMs alone do not constitute equitable evaluators.
@@ 2.2 LLM-based Evaluators
The application of LLM for evaluation purposes has seen significant interest due to its impressive capabilities across various tasks. GPTScore was introduced by Fu et al. (2023) as a method for multi-faceted, tailor-made, and training-independent evaluation. Preliminary investigations by Wang et al. (2023a) into the efficacy of evaluators based on LLM have been conducted. Kocmi and Federmann (2023) have utilised GPT models for assessing machine translation quality. Furthermore, liu-etal-2023-g developed G-EVAL, leveraging GPT-4 to evaluate across multiple tasks such as dialogue response generation, text summarization, data-to-text generation, and machine translation. Despite these advancements, the use of LLM-based metrics to assess adversarial negative responses within open-domain dialogue scenarios remains unexplored.
@@ 3.3 Evaluation Process
After training the SLM, we proceed to the evaluation phase. We first follow G-EVAL liu-etal-2023-g and prompt LLM to evaluate the given context and response. The prompt could be found in  A.1. Therefore we obtain a score referred to as $\mathrm{Score_{LLM}}$. Subsequently, SLM is employed to encode both the context and the response. This step involves segregating the response encoding into two distinct embeddings—robust and non-robust. The mathematical representation of this process is given by
@@ 4.2 Baselines
We select a range of widely used baseline metrics. The word-ovelap and embedding-based metrics includs BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005), Embedding-Average Wieting et al. (2016), Vector-Extrema Forgues et al. (2014), BERTScore Zhang et al. (2020), Unieval Zhong et al. (2022), and BARTScore Yuan et al. (2021). For the LLM-based metrics, we select recently proposed baselines, including G-EVAL liu-etal-2023-g and another LLM-based metric proposed by Chiang and yi Lee (2023) which we denote LLM-Chiang, as well as Gemini.
@@ 4.4.1 Dialogue Classification Task
As Table 1 shows, SLM employs both probability and distance measures, achieves superior performance with an accuracy of 91.05%, which exceeds the 88.91% accuracy achieved by GPT-4. Experimental results show that the SLM particularly excels in the classification of positive responses, achieving the highest accuracy of 91.83%. Conversely, GPT-4 exhibits the strongest performance in classifying adversarial negative responses, achieving 97.40% accuracy. The performance of Gemini surpasses that of GPT-3.5, yet remains inferior to that of GPT-4. These results indicate that while LLMs, such as GPT-4, possess robust capabilities in recognising adversarial negative responses, they still struggle with the one-to-many problem, especially when dealing with semantically diverse positive responses. On the other hand, SLM with the integration of distance and probability metrics shows exceptional performance in recognising open-domain positive responses. Based on these insights, we propose a hybrid approach that combines the SLM, a task-specific model, with LLMs for comprehensive dialogue evaluation. This approach is designed to leverage the distinct strengths of both types of models to enhance the overall performance in open-domain dialogevaluations.
[b]
DailyDialog++ PersonaChat TopicalChat
Metrics Pearson’s $\rho$ Spearman’s $\tau$ Pearson’s $\rho$ Spearman’s $\tau$ Pearson’s $\rho$ Spearman’s $\tau$
BLEU-1 0.303 (<0.0001) 0.242 (<0.0001) 0.331 (<0.0001) 0.276 (<0.0001) 0.352 (<0.0001) 0.330 (<0.0001)
BLEU-2 0.317 (<0.0001) 0.270 (<0.0001) 0.311 (<0.0001) 0.261 (<0.0001) 0.279 (<0.0001) 0.252 (<0.0001)
BLEU-3 0.256 (<0.0001) 0.251 (<0.0001) 0.251 (<0.0001) 0.239 (<0.0001) 0.179 (<0.0001) 0.189 (<0.0001)
BLEU-4 0.234 (<0.0001) 0.247 (<0.0001) 0.214 (<0.0001) 0.217 (<0.0001) 0.247 (<0.0001) 0.154 (<0.0001)
ROUGE-1 0.303 (<0.0001) 0.270 (<0.0001) 0.304 (<0.0001) 0.250 (<0.0001) 0.374 (<0.0001) 0.356 (<0.0001)
ROUGE-2 0.178 (<0.0001) 0.133 (<0.0001) 0.142 (<0.0001) 0.126 (<0.0001) 0.173 (<0.0001) 0.149 (<0.0001)
ROUGE-L 0.305 (<0.0001) 0.271 (<0.0001) 0.294 (<0.0001) 0.244 (<0.0001) 0.279 (<0.0001) 0.259 (<0.0001)
METEOR 0.196 (<0.0001) 0.142 (<0.0001) 0.250 (<0.0001) 0.200 (<0.0001) 0.293 (<0.0001) 0.278 (<0.0001)
Embedding
Extrema 0.382 (<0.0001) 0.367 (<0.0001) 0.330 (<0.0001) 0.299 (<0.0001) 0.389 (<0.0001) 0.244 (<0.0001)
Greedy 0.322 (<0.0001) 0.272 (<0.0001) 0.328 (<0.0001) 0.290 (<0.0001) 0.272 (<0.0001) 0.248 (<0.0001)
Average 0.186 (<0.0001) 0.208 (<0.0001) 0.239 (<0.0001) 0.249 (<0.0001) 0.254 (<0.0001) 0.261 (<0.0001)
BERTScore 0.419 (<0.0001) 0.381 (<0.0001) 0.465 (<0.0001) 0.412 (<0.0001) 0.525 (<0.0001) 0.485 (<0.0001)
BARTSCORE 0.305 (<0.0001) 0.264 (<0.0001) 0.466 (<0.0001) 0.442 (<0.0001) 0.470 (<0.0001) 0.427 (<0.0001)
Unieval 0.103 (<0.0001) 0.069 (<0.0001) 0.192 (<0.0001) 0.131 (<0.0001) 0.357 (<0.0001) 0.290 (<0.0001)
G-EVAL (GPT-3.5) 0.447 (<0.0001) 0.414 (<0.0001) 0.516 (<0.0001) 0.562 (<0.0001) 0.642 (<0.0001) 0.668 (<0.0001)
G-EVAL (GPT-4) 0.641 (<0.0001) 0.570 (<0.0001) 0.597 (<0.0001) 0.663 (<0.0001) 0.696 (<0.0001) 0.729 (<0.0001)
LLM-Chiang (GPT-3.5) 0.632 (<0.0001) 0.583 (<0.0001) 0.599 (<0.0001) 0.670 (<0.0001) 0.667 (<0.0001) 0.661 (<0.0001)
LLM-Chiang (GPT-4) 0.723 (<0.0001) 0.701 (<0.0001) 0.701(<0.0001) 0.675 (<0.0001) 0.678 (<0.0001) 0.709 (<0.0001)
Ours (SLM-Prob-only) 0.677 (<0.0001) 0.632 (<0.0001) 0.638 (<0.0001) 0.692 (<0.0001) 0.691 (<0.0001) 0.689 (<0.0001)
Ours (SLM-Dis-only) 0.652 (<0.0001) 0.600 (<0.0001) 0.668 (<0.0001) 0.692 (<0.0001) 0.642 (<0.0001) 0.673 (<0.0001)
Ours (SLM-Dis and Prob) 0.728 (<0.0001) 0.666 (<0.0001) 0.711 (<0.0001) 0.707 (<0.0001) 0.741 (<0.0001) 0.727 (<0.0001)
Ours (LLMs-only (GPT-3.5)) 0.518 (<0.0001) 0.472 (<0.0001) 0.506 (<0.0001) 0.617 (<0.0001) 0.590 (<0.0001) 0.669 (<0.0001)
SLIDE (GPT-3.5)) 0.727 (<0.0001) 0.674 (<0.0001) 0.692 (<0.0001) 0.695 (<0.0001) 0.696 (<0.0001) 0.718 (<0.0001)
Ours (LLMs-only (Gemini)) 0.660 (<0.0001) 0.614 (<0.0001) 0.719 (<0.0001) 0.688 (<0.0001) 0.701 (<0.0001) 0.699 (<0.0001)
SLIDE (Gemini) 0.760 (<0.0001) 0.689 (<0.0001) 0.701 (<0.0001) 0.705 (<0.0001) 0.713 (<0.0001) 0.712 (<0.0001)
Ours (LLMs-only (GPT-4)) 0.709 (<0.0001) 0.666 (<0.0001) 0.737 (<0.0001) 0.729 (<0.0001) 0.745 (<0.0001) 0.736 (<0.0001)
SLIDE (GPT-4) 0.773 (<0.0001) 0.704 (<0.0001) 0.713 (<0.0001) 0.709 (<0.0001) 0.711 (<0.0001) 0.722 (<0.0001)
@@ 4.4.2 Dialogue Evaluation
Figure 2: T-SNE visualisation of the sentence representation of context and responses. The left panel, labeled Normal, illustrates the vectors prior to disentanglement, whereas the right panel, labeled Disentangled, displays the post-disentanglement outcomes. This demonstrates the convergence of negative responses towards the context following disentanglement.
4.5 Example Visualisation
To demonstrate the effectiveness of disentanglement, we selected a subset of examples from the DailyDialog++ dataset and employed T-SNE for visualisation, as depicted in Figure 2. Each example includes a context, five positive responses, and five adversarial negative responses. The vector representations prior to disentanglement are illustrated in the left plot, whereas those post-disentanglement are shown on the right. Post-disentanglement visualisations reveal a more distinct clustering of responses in relation to their contexts. By utilising these refined representations to train a classifier, we observed an improvement in classification accuracy on the test set of DailyDialog++ dataset, from 87% to 88.19%. Additional examples are available in Appendix A.5.
@@ 4.5 Example Visualisation
FS: We have another traditional holiday-the Dragon Boat Festival.
SS: When is it?
FS: It falls on the fifth day of the fifth lunar month.
@@
FS: We have been over this a hundred times! We are not getting a pet!
SS: Why not? Come on! Just a cute little puppy or a kitty!
FS: Who is going to look after a dog or a cat?
@@
2.00 2.79 2.40
Table 3: Samples from the DailyDialog++ dataset. "FS" is the First Speaker and "SS" is the Second Speaker
4.6 Case Study
5 Conclusion
@@ A.4 Case Study
FS: We have been over this a hundred times! We are not getting a pet!
SS: Why not? Come on! Just a cute little puppy or a kitty!
FS: Who is going to look after a dog or a cat?
@@
FS: Why not? We’re supposed to meet him there.
SS: Why doesn’t he meet us outside?
FS: Why should he do that? It isn’t illegal for us to go in.
@@
5.00 5.00 4.80
Table 4: Samples from the DailyDialog++ dataset. "FS" is the First Speaker and "SS" is the Second Speaker
A.5 Example Visualisation